OpenStudy (aravindg):

Find the value of a for which the function f(x)=asin x+(1/3)sin 3x has an extremum at x=pi/3

OpenStudy (aravindg):

@Mani_Jha , @apoorvk

OpenStudy (apoorvk):

that means at that point, i.e. pi/3 derivative of f(x) is zero. so dy/dx = acosx + cos3x = 0 (at x=pi/3) so, a/2 + (-1) = 0 or, a =2 Samjhe bhole?

OpenStudy (experimentx):

2/sqrt(3) ??? i m not sure ...

OpenStudy (aravindg):

?/

OpenStudy (aravindg):

is der anything wrong with apoorv's wrking according to u @experimentX

OpenStudy (experimentx):

looks like i got right value of cos 30

OpenStudy (aravindg):

ya cos 30 is root3/2

OpenStudy (apoorvk):

pi/3 = 60 degrees.

OpenStudy (experimentx):

Oo .. sorry, my mistake.

OpenStudy (aravindg):

hehe u made a booboo

OpenStudy (aravindg):

tht was simply gr8 @apoorv

OpenStudy (experimentx):

great guys!!

OpenStudy (apoorvk):

oh that wasn't great at all really. :)